Give your answer in scientific notation, rounded to three significant figures.

A sinusoidal EM wave has an electric field amplitude Em = 29.2 mV/m.

What is the intensity and average energy density?

3.77 × 10 J/m^3 (average energy density- figured this part out, don't know how to do the intensity)
Answer #1

here,

Em = 29.2 mV /m = 0.0292 V/m

a)

the intensity of wave, S = Em^2 /( 2*u * c)

s = 0.0292^2 /( 2* 1.26 * 10^-6 * 3 * 10^8)

s = 1.13 * 10^-6 W/m^2

b)

the electric energy density , U = 0.5 * e0 * Em^2

U = 0.5 * 8.85 * 10^-12 * ( 0.0292^2)

U = 3.77 * 10^-15 J/m^3